OpenStudy (anonymous):

Simplify: (7y^2-y+2) + (-6y^2+3y+4)

OpenStudy (earthcitizen):

y^2+2 y+6

OpenStudy (earthcitizen):

collect like terms (7y^2-y+2) + (-6y^2+3y+4) = y^2+2 y+6

OpenStudy (anonymous):

i would begin this problem by distributing anything that is necesary. the only thing that must be distrubted is the + in the middle you would get 7y^2-y+2-6y^2+3y+4 and then you would combine all variables that are exactly the same (x^2,x , and constant) 7y^2-6y^2 = y^2 -y+3y=2y 2+4=6 this gives you y^2+2y+6

OpenStudy (anonymous):

this is simplify